​The probability that a person in the United States has type B blood is 12%.


The probability that a person in the United States has type B blood is 12%. Three unrelated people in the United States are selected at random Complete parts (a) through (d). 

(a) Find the probability that all three have type B blood 


The probability that all three have type B blood is _______ 
Answer #1

Given that

probability of person having B+ in United States will be = 0.12

Probability of having B+ for all 3 selected randomly will be = 0.12*0.12*0.12=0.001728

We have to multiply probability of 3 person because it will be independent event